# Oxygen Vacancies in Niobium Pentoxide as a Source of Two-Level System Losses in Superconducting Niobium

### v3

@@ II Experimental Method
The second resonator, TE1RI003, was used to isolate losses to the dynamics of the oxide itself. After its baseline test post electropolishing, the cavity underwent in situ low temperature baking at 200C for 1 hour followed by further testing. The cavity was then vented to clean room air and left to oxidize for a full week. This drove the re-formation of the native niobium oxide but this time with a larger concentration of subsurface O interstitial present compared to the baseline measurement post EP. Once complete, TE1RI003 was once again re-evacuated and tested.
@@ III.1 Resonator RF measurements
The upper panel in Fig. 1 presents data acquired on resonator TE1AES019 after an initial 90$\times$ 384 hour in situ vacuum bake followed by sequential in situ vacuum bakes at various temperatures and durations. Shown also for reference is data from Romanenko and Schuster taken on an EP resonator which contains a full 5 nm native wet-grown oxide and serves as our baseline test [9]. For all tests, we observe the expected saturation of the quality factor at electric fields $<$ 0.01 MV/m (i.e. photon numbers $<$ 2E18) that is characteristic of two-level systems [9, 10, 6]. We find that the 90$\times$ 384 hours $+$ 200$\times$ 1 hour in situ vacuum baking treatment produces a resonator with a low field quality factor of 9E9, roughly three times lower than the baseline test. An additional 200$\times$ 5 hour in situ vacuum bake produces identical results at fields below 0.1 MV/m. After a second round of 200$\times$ 5 hour vacuum bake, a slight improvement in $Q_{0}$ up to 1.2E10 occurs. For the final treatment, we subjected the resonator to a 340$\times$ 5 hour vacuum baking treatment which is known to fully dissolve the native Nb2O5 layer [10, 40]. The resulting measurements display a $Q_{0}$ at low fields of greater than 7E10, the highest achieved in this study, and in line with previous work [10, 40]. Overall, these tests show a non-monotonic dependence of resonator $Q_{0}$ with bake duration.
@@
$\frac{1}{Q_{0}}=\frac{F\delta_{0}}{\big{(}1+\big{(}\frac{E_{acc}}{E_{c}}\big{)}^{2}\big{)}^{\beta}}+\frac{1}{Q_{qp}}$ (2)
where $\beta$ is a fitting parameter, $E_{c}$ is the characteristic saturation field, $\delta_{0}$ is the dielectric loss, and $Q_{qp}$ is the quasi-particle contribution to RF loss [13, 9]. Fig. 2 plots the product of the filling factor and dielectric loss tangent for each curve presented in Fig. 1. We choose to present this convolved product as it is agnostic to the precise values of the filling factor and loss tangent and gives a qualitative measure of the TLS contribution to RF loss. Resonator TE1AES019 shown in the upper panel of Fig. 2 displays a non-monotonic dependence of the TLS-driven losses with in situ vacuum baking, achieving as low as 3.63E-12 after the final round of treatment, consistent with previous results [10]. TE1RI003 presented in the middle panel of Fig. 2 further suggests that in situ vacuum baking increases TLS-driven RF losses while reoxidation in dry clean room air for one week returns them to baseline EP values. The results of TE1AES021 in the lower panel further support this observation and suggest that a fully oxidized surface, whether dry or wet grown, contains less two-level system loss than a partially dissolved oxide.
Figure 2: Fitted $F\delta_{0}$ values and fit error of a) TE1AES019, b) TE1RI003, and c) TE1AES021 using the data presented in Fig. 1. The 90$\times$ 384 hours in situ vacuum bake that precedes the sequential in situ vacuum baking treatments for resonator TE1AES019 has been omitted from the horizontal axis for brevity. Most error bars are smaller than the symbol size.
We mention that the characteristic saturation field extracted from Fig. 1 remains roughly constant, as evidenced by the increase in $Q_{0}$ at 0.1-0.2 MV/m for all investigated cavities and temperatures. This supports the theory that a single loss mechanism remains dominant throughout these studies: two-level systems located within the oxide.

### v4

@@ II Experimental Method
The second resonator, TE1RI003, was used to isolate losses to the dynamics of the oxide itself. After its baseline test post electropolishing, the cavity underwent in situ low temperature baking at 200C for 1 hour followed by further testing. The cavity was then vented to clean room air and left to oxidize for a full week. This drove the re-formation of the native niobium oxide but this time with a larger concentration of subsurface oxygen (O) interstitial present compared to the baseline measurement post EP. Once complete, TE1RI003 was once again re-evacuated and tested.
@@ III.1 Resonator RF measurements
The upper panel in Fig. 1 presents data acquired on resonator TE1AES019 after an initial 90$\times$ 384 hour in situ vacuum bake followed by sequential in situ vacuum bakes at various temperatures and durations. Shown also for reference is data from Romanenko and Schuster taken on an EP resonator which contains a full 5 nm native wet-grown oxide and serves as our baseline test [9]. For all tests, we observe the expected saturation of the quality factor at electric fields $<$ 0.01 MV/m (i.e. photon numbers $<$ 2E18) that is characteristic of two-level systems [9, 10, 6]. We find that the 90$\times$ 384 hours $+$ 200$\times$ 1 hour in situ vacuum baking treatment produces a resonator with a low field quality factor of 9E9, roughly three times lower than the baseline test. An additional 200$\times$ 5 hour in situ vacuum bake produces identical results at fields below 0.1 MV/m. After the third round of treatment of 200$\times$ 5 hour vacuum bake, a slight improvement in $Q_{0}$ up to 1.2E10 occurs. For the final treatment, we subjected the resonator to a 340$\times$ 5 hour vacuum baking treatment which is known to fully dissolve the native Nb2O5 layer [10, 38]. The resulting measurements display a $Q_{0}$ at low fields of greater than 7E10, the highest achieved in this study, and in line with previous work [10, 38]. Overall, these tests show a non-monotonic dependence of resonator $Q_{0}$ with bake duration.
@@
$\frac{1}{Q_{0}}=\frac{F\delta_{TLS}(T)}{\big{(}1+\big{(}\frac{E_{acc}}{E_{c}}\big{)}^{2}\big{)}^{\beta}}+\frac{1}{Q_{qp}}$ (2)
where $\beta$ is a fitting parameter, $E_{c}$ is the characteristic saturation field, $\delta_{TLS}$(T) is the temperature dependent TLS dielectric loss tangent, and $Q_{qp}$ is the quasi-particle contribution to RF loss [39, 9]. Fig. 2 plots the product of the filling factor and dielectric loss tangent for each curve presented in Fig. 1. We choose to present this convolved product as it is agnostic to the precise values of the filling factor and loss tangent and gives a qualitative measure of the TLS contribution to RF loss. Resonator TE1AES019 shown in the upper panel of Fig. 2 displays a non-monotonic dependence of the TLS-driven losses with in situ vacuum baking, achieving as low as 3.63E-12 after the final round of treatment, consistent with previous results [10]. TE1RI003 presented in the middle panel of Fig. 2 further suggests that in situ vacuum baking increases TLS-driven RF losses while reoxidation in dry clean room air for one week returns them to baseline EP values. The results of TE1AES021 in the lower panel further support this observation and suggest that a fully oxidized surface, whether dry or wet grown, contains less two-level system loss than a partially dissolved oxide.
Figure 2: Fitted $F\delta_{TLS}(T)$ values and fit error of a) TE1AES019, b) TE1RI003, and c) TE1AES021 using the data presented in Fig. 1 for 1.4 K $<T<$ 1.6 K. The 90$\times$ 384 hours in situ vacuum bake that precedes the sequential in situ vacuum baking treatments for resonator TE1AES019 has been omitted from the horizontal axis for brevity. Most error bars are smaller than the symbol size.
We mention that the characteristic saturation field extracted from Fig. 1 remains roughly constant, as evidenced by the increase in $Q_{0}$ at 0.1-0.2 MV/m for all investigated cavities and temperatures. This supports the theory that a single loss mechanism remains dominant throughout these studies: two-level systems located within the oxide. However, such a critical field is considerably higher than the values of $\sim$1 V/m observed at mK in 2-D devices [40, 41]. We believe that this is due to two factors. First, the large loss tangent of the pentoxide [10] may require a large electric field (i.e. more photons) to saturate the TLS. Moreover, due to the inhomogeneous electric field distribution within the cavities, when the reported on-axis field is $E_{acc}\approx$ 0.1 MV/m, surface TLS interact with electric fields which vary from $\sim$0 - 8e5 V/m, as depicted in Fig. 3. Second, the data reported in Fig. 1 was obtained at comparatively higher temperatures of 1.4 K-1.6 K; it has been theoretically and experimentally shown that $E_{c}$ increases with increasing temperature [42]. As a result, the values of $E_{c}$ are in line with what might be expected from the conventional theory of TLS.
Figure 3: Electric field distribution with $E_{acc}$approx 0.1 MV/m.